OpenStudy (anonymous):

In the system of x+2y=9 and x-y=6. What would be substituted for x in the second equation? A.9-x B.1/2x+9 C.9-2y D.y-9/2 I got D. Am I right?

OpenStudy (anonymous):

guess what i'm gonna say!!

OpenStudy (anonymous):

I got y=-1/2x+9/2

OpenStudy (anonymous):

what are you going to say?

OpenStudy (anonymous):

What are you gonna say?

OpenStudy (anonymous):

Wait so Your saying it's 'B'?

OpenStudy (anonymous):

nope its C

OpenStudy (anonymous):

Can you explain that please.

OpenStudy (anonymous):

sorry about that i was trying to find out what why was equal to.

OpenStudy (anonymous):

so all you have to do is change the equation x+2y=9 to what x=. subtract 2y to the other side

OpenStudy (anonymous):

x+2y=9 -2y -2y x = 9 - 2y ------------ x - y = 6 (9-2y) - y = 6

OpenStudy (anonymous):

you get -2y+9 which is the same as 9-2y

OpenStudy (anonymous):

ooh..

OpenStudy (anonymous):

therefore the answer is C

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

hint: Step 1) type the two equations in wolfram alpha separated by a comma Step 2) ??? Step 3) Sing! Step 4) ANSWER!!!

OpenStudy (anonymous):

Hahaha

OpenStudy (anonymous):

thats stupid

OpenStudy (anonymous):

you'll thank me by the time you get to Calculus 3! XD

OpenStudy (anonymous):

ok ok , so can you explain this one too. Solve this system -5x+y=-3 3x-8y=24 (0,-3) (3,0) (0,3) (-3,0)

OpenStudy (anonymous):

use the hint I gave you!

OpenStudy (anonymous):

How will that help me?